A man of mass 50kg ascends a flight of stairs 5m high in 5 seconds. If accelerator due to gravity is 10m-2. Calculate the power expended.

1
Expert's answer
2022-01-07T09:29:39-0500

Given:

"m=50\:\rm kg"

"h=5\:\rm m"

"t=5\:\rm s"


The power

"P=\frac{W}{t}=\frac{mgh}{t}"

"P=\frac{50*10*5}{5}=500\rm\: W"